# An Algebraic Roadmap of Particle Theories Part I: General construction

### v1

@@ VI Conclusion
In what may be viewed as a consistency condition, we introduced what we call the multivector condition. This constraint dictates that minimal left ideal fermions transform in accordance with the multivectors that comprise them. As a result, for internal symmetries, we find that the adjoint, vector, and spinor representations now each transform infinitesimally under the commutator. The minimal left ideal spinors that we constructed constitute a special case where the commutator and left multiplication coincide. The condition is algebraically equivalent to, but conceptually distinct from, constraints of earlier authors fixing the volume form of a $\Lambda\mathbb{C}^{5}$ spinor.
@@
The findings of Baez and Huerta, [1], and Furey and Hughes, [2], [3], are fused into a single particle roadmap. A total of nine particle models are shown to be interlinked via the multivector condition, and via constraints of invariance under generalized division algebraic reflections. Six of these models are familiar to particle physicists, namely the “SO(10)” model, the Georgi-Glashow model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model pre-Higgs mechanism, and the Standard Model post-Higgs mechanism.
@@ Acknowledgements.
These manuscripts have benefitted from numerous discussions with Beth Romano. The author is furthermore grateful for feedback and encouragement from John Baez, Sukṛti Bansal, John Barrett, Latham Boyle, Hilary Carteret, Mia Hughes, Kaushlendra Kumar, Agostino Patella, Shadi Tahvildar-Zadeh, Carlos Tamarit, Andreas Trautner, and Jorge Zanelli. This work was graciously supported by the VW Stiftung Freigeist Fellowship, and Humboldt-Universität zu Berlin.

### v2

@@ VII Conclusion
In what may be viewed as a consistency condition, we introduced what we call the multivector condition. This constraint dictates that minimal left ideal fermions transform in accordance with the multivectors that comprise them. As a result, for internal symmetries, we find that the adjoint, vector, and spinor representations now each transform infinitesimally under the commutator. The minimal left ideal spinors that we constructed constitute a special case where the commutator and left multiplication coincide.
@@
The findings of Baez and Huerta, [1], and Furey and Hughes, [2], [3], are reconstructed, extended, and amalgamated into a single particle roadmap. A total of nine particle models are shown to be interlinked via the multivector condition, and via constraints of invariance under generalized division algebraic reflections. Six of these models are familiar to particle physicists, namely the “SO(10)” model, the Georgi-Glashow model, the Pati-Salam model, the Left-Right Symmetric model, the Standard Model pre-Higgs mechanism, and the Standard Model post-Higgs mechanism.
@@ Acknowledgements.
These manuscripts have benefitted from numerous discussions with Beth Romano. The author is furthermore grateful for feedback and encouragement from John Baez, Sukruti Bansal, John Barrett, Latham Boyle, Hilary Carteret, Mia Hughes, Kaushal Kumar, Agostino Patella, Shadi Tahvildar-Zadeh, Carlos Tamarit, Andreas Trautner, and Jorge Zanelli. This work was graciously supported by the VW Stiftung Freigeist Fellowship, and Humboldt-Universität zu Berlin.